If the first three terms of an arithmetic sequence are (2x+1,4x+2,10x-1,...), what is the value of x? Please can someone explain ? I think the answer is 1 but I don't really understand this question.
Answer by mayank(15)


Ans.The answer to the above question is x=1. It is because as the given three terms are in arithmetic sequence, the difference betweent two consecutive terms remains constant and same.Thus 2nd term-1st term= 3rd term-2nd term.Therefore,
4x+2-(2x+1)=10x-1-(4x+2)
thus,4x+2-2x-1=10x-1-4x-2
thus 2x+1=6x-3
thus 4x=4
and this leads to x=1.